OpenStudy (anonymous):

Alycia opens her piggy bank to find only dimes and quarters inside. She counts the coins and finds that she has 47 coins totaling $7.70. How many dimes were in the bank, and how many quarters were in there? (use linear equation to solve)

OpenStudy (anonymous):

From the problem statement: {10 d + 25 q = 770, d + q = 47} Solve for d and q. {d = 27, q = 20} (27*0.10) + (20*0.25) = $7.70

OpenStudy (anonymous):

Thanks!